OpenStudy (anonymous):

Find three positive numbers whose sum is 100 and whose product is a maximum.

OpenStudy (anonymous):

@amistre64 How do I do this without the use of Lagrange Multipliers?

OpenStudy (cwrw238):

hhmmm i must admit i don't know what lagrange multipliers are but intuition tells me they are 32, 33 and 34

OpenStudy (cwrw238):

assuming we are talking whole numbers

OpenStudy (anonymous):

When they dont have to be whole numbers they are 100/3 = 33.333 But I don't know how to do them with fx = 0 fy = 0 -that method

OpenStudy (phi):

how about this: x+y+z = 100 ---> z = 100-x-y f(x,y,z) = x y z f(x,y)= x y (100 - x -y) = 100xy -x^2 y - xy^2 fx= 100y -2 xy - y^2 fy = 100x - x^2 - 2 x y set fx =0 and fy =0 solve for x (or y)

OpenStudy (anonymous):

Ah thanks! I didn't think of getting z and plugging that into the xyz equation! Thanks!!